### v3

@@ 3.1 Problem Identification
In our research, we sought to assess the ergonomic risks that students face due to prolonged sitting at their desks and the potential development of Work-related MSDs (WMSDs). We developed a comprehensive self-reported questionnaire based on the Nordic Musculoskeletal Questionnaire (NMQ), which has been extensively used and validated in previous studies. The questionnaire was designed to gather information on the frequency and severity of musculoskeletal symptoms experienced by teachers, focusing on areas such as the neck, shoulders, back, and wrists, which are commonly affected by prolonged sitting and poor ergonomics. Participants were asked to report the frequency of musculoskeletal symptoms they experienced, classifying them as “constantly” (most of the time during the day), “occasionally” (two to four times a month), or “frequently” (more than four times a month). We also asked the participants whether they had experienced musculoskeletal pain in the past 12 months. This allowed us to assess the prevalence of musculoskeletal pain among teachers and to identify any data patterns or trends. Approximately 78.42% of students reported experiencing MSD-related pains. In this group, 77.67% were male, and 81.25% were female. We collected demographic information, such as age, gender, and level of study, to better understand how these factors might influence the prevalence and severity of musculoskeletal symptoms. The questionnaire was administered to a sample of KUET students. The collected data were analyzed to identify any significant associations between ergonomic factors, such as prolonged sitting and poor posture, and the development of WMSDs. The reliability test for the NMQ survey yielded a Cronbach’s alpha value of 0.77, indicating an “acceptable” level of internal consistency. This value falls within the 95% confidence interval of 0.733 and 0.803, suggesting that if the survey were repeated with different samples, the actual value of Cronbach’s alpha would likely fall within this range 95% of the time.
@@ 3.2 Sample Collection
A subset of students was chosen from two departments at the Khulna University of Engineering & Technology (KUET), Khulna, Bangladesh. Equation 1 was used to establish the necessary sample size.
@@
where, $n$ represents the sample size or population size, $N$ denotes the total population, and $e$ refers to the desired precision level set at 5% with a 95% confidence level.
According to KUET’s official database, there are currently 5,240 students enrolled, comprising 4,318 males and 922 females. This yielded a male-to-female ratio of approximately 4.68:1. According to Equation 1, a minimum of 372 or more data points should be collected. This study gathered anthropometric data from 300 males and 80 females, representing up to 7,000 students. All the students’ ages ranged from 17 to 26 years. The Director of Research & Extension, KUET, approved the experimental procedures, which were conducted in accordance with the principles stated in the Declaration of Helsinki. Participation in the survey was voluntary, and participants had the right to withdraw at any time without penalties. Confidentiality of participant responses was strictly maintained, and all data collected were anonymized and stored securely. The dataset, available for further exploration, can be found in “Anthropometric Data of KUET students(Jahin and Saha,, 2024).
Table 1: Demographic distribution of study population by gender and academic level
Study level Male Female
1st year student 70 20
2nd year student 70 20
3rd year student 70 20
4th year student 90 20
Total 300 80
@@ 3.3 Anthropometric Measurements
The 11 anthropometric measurements described in Table 2 were obtained for this study (see Figure 2). These dimensions were measured using definitions employed in earlier studies (Kroemer et al.,, 1997). A chair and a Harpenden Anthropometer were used to determine the seating dimensions, and sitting assistance was used to support participants in maintaining a straight posture with their feet flat on the floor. Before data collection, rigorous calibration of all equipment was conducted in accordance with the established standards to maintain the accuracy and precision of the measurements. Throughout the trial, the participants were instructed to maintain a straight posture on a flat-surfaced seat, with their knees positioned at a $90^{\circ}$ angle. Standardized posture is crucial for the consistency and comparability of anthropometric data. Participants were briefed about the study objectives, and precise postures were maintained during anthropometric measurements. All measurements were documented in millimeters (mm).
(a) Right-side view
(b) Backside view
Figure 2: Anthropometry of the participants, (a) right-side view, (b) backside view.
Table 2: Anthropometric measurements used in this study (International Organization for Standardization,, 2017)
Anthropometric dimensions Acronym Description
Popliteal Height PH
Vertical distance from the floor to the popliteal angle, located on the lower surface of the knee, while keeping the feet flat on the floor.
Sitting Elbow Height SEH
Distance from the seat surface to the underside of the elbow, with the forearm horizontal and the upper arm hanging freely downwards.
Buttock Popliteal Length BPL
Distance from the back of the seat to the back of the knee, where the backs of the lower legs meet the underside of the thighs.
Buttock Knee Length BKL
Horizontal distance from the back of the buttocks to the back of the knee.
Hip Breadth HB
The horizontal measurement taken across the hips while the participant is sitting and keeping their knees together (maximum width).
Sitting Shoulder Height SSH
Vertical distance from the seat surface to the acromion (bony shoulder point).
Sitting Elbow Height SEB
Horizontal distance across the lateral surfaces of the elbows.
Thigh Thickness TT
Vertical distance from the seat surface to the top of the thigh, where the uncompressed soft tissue of the thigh meets the abdomen.
Arm Length AL
Vertical distance from the scapula’s lowest point (inferior angle) to the participant’s sitting surface.
Elbow Fingertip Length EFL
Horizontal distance from the back of the elbow to the tip of the middle finger in a standard sitting position.
Subscapular Height SCH
Vertical distance from the scapula’s lowest point (inferior angle) to the participant’s sitting surface.
@@ 3.4 Furniture Dimensions
The dimensions used for designing the furniture are shown in Figure 3 with identities. This study evaluated two types of university furniture used in computer labs at KUET. One was a ‘table with adjustable chairs,’ and the other was a ‘table with non-adjustable chairs,’ both commonly used in computer labs at KUET. Figure 3 depicts the computer lab furniture design parameters, whereas Figure 4 shows the two furniture sets. As such, the study measured 11 design parameters for the two types of existing computer furniture, as described in Table 3.
@@
Table 3: Parameters used for furniture design in this study
Furniture dimension Acronym Description
Seat Height SH Vertical distance from the floor to the top of the seat surface.
Seat Width SW Horizontal distance between the lateral edges of the seat.
Seat Depth SD Horizontal distance from the front edge to the back edge of the seat surface.
Backrest Height BH Vertical distance from the seat surface’s top edge to the backrest’s top edge.
Backrest Width BW Horizontal distance between the lateral edges of the backrest.
Seat to Table Height STH Vertical distance from the seat surface’s top edge to the table surface’s top edge.
Table Depth TD Horizontal distance from the front edge to the back edge of the table surface.
Table Length TL The maximum horizontal distance between the lateral edges of the table surface
Table Height TH Vertical distance from the floor to the top edge of the table surface.
Underneath Table Height UTH Vertical distance from the lowest point below the table to the floor’s surface.
Upper Edge of Backrest UEB Vertical distance from the backrest’s top edge to the headrest’s top edge.
@@ 3.5 Furniture Dimensions and Anthropometric Data Mismatch
Considering the students’ anthropometric measurements when designing and evaluating computer lab furniture is essential. This involves using relevant anthropometric estimations and ergonomic standards to determine the acceptable range for each furniture measurement.
3.5.1 PH against SH
Popliteal height (PH) and seat height (SH) are typically connected. SH should provide knee flexion to align the lower legs with the body’s craniocaudal axis at a $30^{\circ}$ angle (JFM et al.,, 2003). According to Parcells et al., (1999), PH should be higher than SH. This means that the shin-thigh position should be between $95^{\circ}$ and $120^{\circ}$, and the lower leg should create an angle with a vertical axis of $5^{\circ}$ to $30^{\circ}$. While an unduly low SH can strain the ischial tuberosities, an extremely high SH can increase the pressure on the bottom of the knee and nerves, decreasing blood circulation (Hoque et al.,, 2014). The SH should be less than 40 mm, or 88% of the PH, to prevent putting excessive pressure on the buttocks (JFM et al.,, 2003). The PH measurement considered the shoe height by adding 30 mm. Therefore, Equation 2 can be used to represent the relationship between PH and SH:
@@ 3.5.1 PH against SH
3.5.2 HB against SW
To allow for both furniture flexibility and essential motions, the seat width (SW) should be broad enough to accommodate people with the widest hip breadth (Evans et al.,, 1988; Sanders and McCormick,, 1993). A proper fit is suggested by research, and the 95th percentile of HB is a good place to start when measuring SW (Evans et al.,, 1988; Occhipinti et al.,, 1993; Oborne,, 1995; Sanders and McCormick,, 1993). The SW should be wider than the HB. Previous research has suggested that the optimal SW should to be within the range of 10% to 30% of HB (Dianat et al.,, 2013; Gouvali and Boudolos,, 2006). Consequently, the link between the HB and SW can be established using Equation 3:
@@ 3.5.2 HB against SW
3.5.3 BPL against SD
There is a strong relationship between SD and BPL, with BPL frequently serving as a guide for SD development and evaluation. The fifth percentile of the BPL should be used as the basis for the seat depth, according to several studies (Hossain and Ahmed,, 2010; Sanders and McCormick,, 1993; Oborne,, 1995; Panagiotopoulou et al.,, 2004). This guarantees that the seat backrest can support the lumbar spine without compressing the popliteal surface. The seat may not support the thigh when the SD is much shorter than the BPL (Castellucci et al.,, 2010). On the other hand, the user might be unable to use the backrest to support their lumbar spine if the SD is larger than the BPL (Milanese and Grimmer,, 2004). Accordingly, when SD is either 95% of BPL or less, there is a mismatch between the SD and BPL (Afzan et al.,, 2012; Altaboli et al.,, 2015; Dianat et al.,, 2013). Equation 4 can be used to express the connection between BPL and SD.
@@ 3.5.3 BPL against SD
3.5.4 SEH against STH
According to several studies, the SEH is a critical factor in SDH determination (Bendix and Bloch,, 1986; Milanese and Grimmer,, 2004; JFM et al.,, 2003). Developing an appropriate TH requires careful consideration of shoulder biomechanics, including shoulder flexion angles between $0^{\circ}$ and 2525$\circ$25 ∘ and abduction angles between $0^{\circ}$ and $20^{\circ}$ (Agha,, 2010; Altaboli et al.,, 2015; Gouvali and Boudolos,, 2006). Furthermore, keeping the arms on the desk significantly reduces strain on the spine. Consequently, according to ref (Poulakakis and Marmaras,, 1998)[50 Pheasant, 51 Poulakakis], the STH should be 3-5 cm higher than the SEH. A synchronized measure was created to verify that the SEH is the most important TH parameter and that the TH height should not increase above 50 mm above the SEH. Equation 5 expresses the link between TH and SEH.
@@ 3.5.4 SEH against STH
3.5.5 SSH against BH
Flexibility of the arms and upper trunk depends on the appropriate BH. It is recommended that the BH value is not higher than that of the scapula [31 Evans]. Therefore, an ideal BH is between 60% and 80% of sitting shoulder height (Agha,, 2010; Dianat et al.,, 2013; Castellucci et al.,, 2015). BH dimensions should also be less than or equal to the top margin of the scapula. This creates a match criterion, and Equation 6 shows how BH and SSH relate to one another.
@@ 3.5.5 SSH against BH
3.5.6 TT against SDC
The SDC should be sufficiently large to provide adequate space for leg movement and for the user to move the chair below the table (García-Acosta and Lange-Morales,, 2007; JFM et al.,, 2003). If the SDC is larger than the TT, it is considered a standard dimension (Parcells et al.,, 1999). However, the ideal dimension of SDC should be 20 mm higher than the knee height (Gouvali and Boudolos,, 2006). Therefore, a match criterion was developed, and the relationship between SDC and TT is shown in Equation 7.
$(TT+20)<SDC$ (7)
3.5.7 PH and TT against UTH
The user should have support for simple access to the seat, leg mobility, and getting off the seat from the Underneath Table Height (UTH) or Under Desktop Height (UDH). In addition, the design must incorporate SH, TT, and an additional 20 mm of space to allow for a shift in leg posture and the lowest standard dimension of UTH (Pérez-Gosende,, 2017). However, according to Gouvali and Boudolos, (2006), the highest dimension of UTH cannot be greater than the difference between the maximum dimension of TH and the table’s thickness, which is 30mm. Consequently, the link between UTH, PH, and TT yields a match equation, as shown in Equation 8.
$(SH+TT+30)\leq UTH\leq(SEH+[(PH+30)\cos(5^{\circ})]+0.1483AL-30)$ (8)
3.5.8 HB against BW
The HB is regarded as a key metric in the BW design process. The dimensions of BW have been considered by several authors (Thariq,, 2010). Taifa and Desai, (2017) has recently suggested that HB is considered a pertinent anthropometric measurement for BW design. Consequently, an equation was created, and Equation 9 illustrates the relationship between HB and BW:
@@ 3.5.8 HB against BW
3.5.9 BKL against TL
The TL is crucial for providing sufficient room for shifting lower body posture or movement. Previous research has suggested that TL should be higher than BKL (Pérez-Gosende,, 2017). Thus, Equation 10 defines the TL minimal limit.
@@ 3.5.9 BKL against TL
3.5.10 SEB against TD
The TD and SEB anthropometric measurements were connected. The partial widths of the AL and SEB should be included in the lower bound of the TD (Pérez-Gosende,, 2017). Sufficient room is needed for a table so that one can enter and leave the seat. TD thus permits elbow abduction at $20^{\circ}$ with a 20 mm margin (Afzan et al.,, 2012; Agha,, 2010; Altaboli et al.,, 2015; Gouvali and Boudolos,, 2006). Equation 11 considers these factors.
@@ 3.5.10 SEB against TD
3.5.11 SCH against UEB
Another important anthropometric criterion for evaluating UEB is SCH (Bendix and Bloch,, 1986). Numerous studies have indicated that the scapula and arm are immobile if the SCH is smaller than the UEB (Castellucci et al.,, 2010; García-Acosta and Lange-Morales,, 2007; Gouvali and Boudolos,, 2006). Equation 12, therefore, depicts the link between SCH and UEB.
@@ 3.6 Level of Compatibility
Some equations have single-directional limitations, while others have dual constraints to assess how well participants’ anthropometric measurements match furniture dimensions. Match and mismatch categories were designed to evaluate compatibility in single-directional connections. Three categories were established for the dual-directional relationships: (a) match (where the anthropometric measurement falls between the limits), (b) low mismatch (where the anthropometric measurement falls below the maximum limit of the relationship), and (c) high mismatch (where the minimum limit of the relationship exceeds the anthropometric measurement).
4 Results and Discussion

### v4

@@ 3.1 Problem Identification
In our study, we sought to assess the ergonomic risks that students face due to prolonged sitting at their desks and the potential development of Work-related MSDs (WMSDs). We developed a comprehensive self-reported questionnaire based on the Nordic Musculoskeletal Questionnaire (NMQ), which has been extensively used and validated in previous studies. The questionnaire was designed to gather information on the frequency and severity of musculoskeletal symptoms experienced by teachers, focusing on areas such as the neck, shoulders, back, and wrists, which are commonly affected by prolonged sitting and poor ergonomics. Participants were asked to report the frequency of musculoskeletal symptoms they experienced, classifying them as “constantly” (most of the time during the day), “occasionally” (two to four times a month), or “frequently” (more than four times a month). We also asked participants whether they had experienced musculoskeletal pain in the past 12 months. This allowed us to assess the prevalence of musculoskeletal pain among teachers and identify any data patterns or trends. Approximately 78.42% of students reported experiencing MSD-related pain. In this group, 77.67% of the patients were male, and 81.25% were female. To better understand how these factors might influence the prevalence and severity of musculoskeletal symptoms, we collected demographic information, such as age, sex, and level of study. The questionnaire was administered to a sample of students of Khulna University of Engineering & Technology (KUET). The collected data were analyzed to identify any significant associations between ergonomic factors, such as prolonged sitting and poor posture, and the development of WMSDs. The reliability test for the NMQ yielded a Cronbach’s alpha value of 0.77, indicating an “acceptable” level of internal consistency. This value falls within the 95% confidence interval of 0.733 and 0.803, suggesting that if the survey were repeated with different samples, the actual value of Cronbach’s alpha would likely fall within this range 95% of the time.
@@ 3.2 Sample Collection
A subset of students was chosen from two departments at KUET, Khulna, Bangladesh. Equation 1 was used to establish the necessary sample size.
@@
where $n$ represents the sample or population size, $N$ denotes the total population, and $e$ refers to the desired precision level set at 5% with a 95% confidence level.
According to KUET’s official database, there are currently 5,240 students enrolled, comprising 4,318 males and 922 females. This yielded a male-to-female ratio of approximately 4.68:1. According to Equation 1, a minimum of 372 or more data points should be collected. This study gathered anthropometric data from 300 males and 80 females, representing up to 7,000 students. The students’ ages ranged from 17 to 26 years. We collected data from June 2023 to September 2023.
All participants provided written consent for the study after being briefed on its aims. The Office of the Director for Research & Extension, KUET, approved the experimental procedures conducted according to the principles stated in the Declaration of Helsinki, bearing the authorization code KUET/DRE/2023/15(6). Participation in the survey was voluntary, and participants had the right to withdraw without penalties. The confidentiality of participants’ responses was strictly maintained, and all data collected were anonymized and stored securely. The dataset, available for further exploration, can be found in “Anthropometric Data of KUET students” (https://dx.doi.org/10.17632/kw7fd465v7.2) [33].
@@ 3.3 Anthropometric Measurements
The 11 anthropometric measurements described in Figure 1 were obtained in this study. These dimensions were measured using definitions employed in earlier studies [34]. A chair and Harpenden Anthropometer were used to determine the seating dimensions, and sitting assistance was used to support participants in maintaining a straight posture with their feet flat on the floor. Before data collection, rigorous calibration of all equipment was conducted in accordance with the established standards to maintain the accuracy and precision of the measurements. Throughout the trial, the participants were instructed to maintain a straight posture on a flat-surfaced seat with their knees positioned at a $90^{\circ}$ angle. Standardized posture is crucial for the consistency and comparability of anthropometric data. Participants were briefed about the study objectives, and precise postures were maintained during anthropometric measurements. All measurements were documented in millimeters (mm). Anthropometric measurements were performed by trained measurers following standardized techniques as outlined by the International Society for the Advancement of Kinanthropometry (ISAK) guidelines. All measurements were performed twice to ensure accuracy and reliability. The average of the two measurements was used for the data analysis.
Figure 1: Anthropometry of the participants, (a) right-side view (on the left), (b) backside view (on the right).
@@ 3.4 Furniture Dimensions
The dimensions used to design the furniture are shown in Figure 2. This study evaluated two types of university furniture sets used in KUET computer laboratories. One was a ‘table with adjustable chairs,’ and the other was a ‘table with non-adjustable chairs,’ both commonly used in computer labs at KUET. Figure 2 depicts the computer lab furniture design parameters, and Figure 3 shows the isometric view of two existing furniture sets. As such, this study measured 11 design parameters of existing computer furniture.
@@ 3.5 Furniture Dimensions and Anthropometric Data Mismatch
Considering the students’ anthropometric measurements when designing and evaluating computer lab furniture is essential. This involves the use of relevant anthropometric estimations and ergonomic standards to determine the acceptable range for each furniture measurement.
Popliteal height (PH) and seat height (SH) are typically connected. The SH should provide knee flexion to align the lower legs with the body’s craniocaudal axis at a $30^{\circ}$ angle [35]. According to [36], PH should be higher than SH. This means that the shin-thigh position should be between $95^{\circ}$ and $120^{\circ}$, and the lower leg should create an angle with a vertical axis of $5^{\circ}$ to $30^{\circ}$. While an unduly low SH can strain the ischial tuberosities, an extremely high SH can increase the pressure on the bottom of the knee and nerves, thereby decreasing blood circulation [16]. The SH should be less than 40 mm, or 88% of the PH, to prevent excessive pressure on the buttocks [35]. The PH measurement considered shoe height by adding 30 mm. Therefore, Equation 2 can be used to represent the relationship between PH and SH:
@@
To allow for both furniture flexibility and essential motions, the seat width (SW) should be broad enough to accommodate people with the widest hip breadth (HB) [37, 38]. A proper fit is suggested by previous research, and the 95th percentile of HB is a good place to start when measuring SW [37, 39, 40, 38]. The SW should be wider than the HB. Previous research has suggested that the optimal SW should be within the range of 10% to 30% of HB [41, 42]. Consequently, the link between the HB and SW can be established using Equation 3:
@@
There is a strong relationship between seat depth (SD) and buttock popliteal length (BPL), with the BPL frequently serving as a guide for SD development and evaluation. The fifth percentile of the BPL should be used as the basis for the seat depth, according to several studies [38, 40, 43]. This guarantees that the seat backrest can support the lumbar spine without compressing the popliteal surface. The seat may not support the thigh when the SD is much shorter than the BPL [44]. However, the user might be unable to use the backrest to support their lumbar spine if the SD is larger than the BPL [45]. Accordingly, when the SD is either 95% of BPL or less, there is a mismatch between the SD and BPL [46, 47, 41]. Equation 4 can be used to express the connection between BPL and SD:
@@
According to several studies, sitting elbow height (SEH) is a critical factor in Seat to Table Height (STH) determination [48, 45, 35]. Developing an appropriate Table Height (TH) requires careful consideration of shoulder biomechanics, including shoulder flexion angles between $0^{\circ}$ and 2525$\circ$25 ∘ and abduction angles between $0^{\circ}$ and $20^{\circ}$ [49, 47, 42]. Furthermore, keeping the arms on the desk significantly reduces strain on the spine. Consequently, according to ref [50][50 Pheasant, 51 Poulakakis], the STH should be 3-5 cm higher than the SEH. A synchronized measure was created to verify that the SEH is the most important TH parameter and that the TH height should not increase above 50 mm above the SEH. Equation 5 expresses the link between TH and SEH:
@@
The flexibility of the arms and upper trunk depends on the appropriate backrest height (BH). It is recommended that the BH value should not be higher than that of the scapula [37]. Therefore, an ideal BH is between 60% and 80% of sitting shoulder height (SSH) [49, 41, 51]. BH dimensions should also be less than or equal to the top margin of the scapula. This creates a match criterion, and Equation 6 shows how BH and SSH relate to each other:
@@
The seat-to-table clearance (STC) should be sufficiently large to provide adequate space for leg movement and for the user to move the chair below the table [52, 35]. If the STC is greater than the thigh thickness (TT), it is considered a standard dimension [36]. However, the ideal dimension of STC should be 20 mm higher than the knee height [42]. Therefore, a match criterion was developed, and the relationship between SDC and TT is shown in Equation 7:
$(TT+20)<STC$ (7)
The user should have support for simple access to the seat, leg mobility, and getting off the seat from under the hood (UTH). In addition, the design must incorporate SH, TT, and an additional 20 mm of space to allow for a shift in leg posture and the lowest standard dimension of UTH [53]. However, according to [42], the highest dimension of UTH cannot be greater than the difference between the maximum dimension of TH and the table thickness, which is 30mm. Consequently, the link between UTH, PH, and TT yields a match equation, as shown in Equation 8:
$(SH+TT+30)\leq UTH\leq(SEH+[(PH+30)\cos(5^{\circ})]\\ +0.1483AL-30)$ (8)
Hip breadth (HB) is a key metric in the backrest width (BW) design process [14]. Taifa and Desai recently suggested that HB is a pertinent anthropometric measurement for BW design [15]. Consequently, an equation was created, and Equation 9 illustrates the relationship between HB and BW:
@@
The table length (TL) is crucial for providing sufficient room for shifting lower body postures or movements. Previous research has suggested that TL should be higher than buttock knee length (BKL) [53]. Thus, Equation 10 defines the TL minimal limit.
@@
The table depth (TD) and sitting elbow breath (SEB) anthropometric measurements were connected. The partial widths of the arm length (AL) and SEB should be included in the lower bound of the table depth (TD) [53]. Sufficient room is needed for a table so that one can enter and leave the seat. TD thus permits elbow abduction at $20^{\circ}$ with a 20 mm margin [46, 49, 47, 42]. Equation 11 considers these factors as follows:
@@
Another important anthropometric criterion for evaluating the upper edge of the backrest (UEB) is the sub-scapular height (SCH) [48]. Numerous studies have indicated that the scapula and arm are immobile if the SCH is smaller than the UEB [44, 52, 42]. Equation 12, therefore, depicts the link between SCH and UEB.
@@ 3.6 Level of Compatibility
Some equations have single-directional limitations, while others have dual constraints to assess how well participants’ anthropometric measurements match the furniture dimensions. Match and mismatch categories were designed to evaluate compatibility in single-directional connections. Three categories were established for dual-directional relationships: (a) match (where the anthropometric measurement falls between the limits), (b) low mismatch (where the anthropometric measurement falls below the maximum limit of the relationship), and (c) high mismatch (where the minimum limit of the relationship exceeds the anthropometric measurement).
3.7 Software and Statistical Tools
IBM SPSS Statistics 26 and Microsoft Excel 2021 were used for statistical analysis. For the graphical 3D prototype design, we used the Autodesk Fusion 360. For visualizations related to anthropometric data distribution and correlation analysis, Python 3.10.
4 Results